Given this problem:
Max Z = $0.3x + $0.90y
Subject to:
2x + 3.2y <= 160
4x + 2.0y <= 240
y <= 40
x, y >= 0


a) Solve for the quantities of x and y which will maximize Z. The x =. The y = .


b) What is the maximum value of Z? The Z =
Answer #1

Given problem is linear programming , and no method precribe to solve hence solving problem by Graphical method

Max Z =$0.30 x +$0.90y

Subject to Contraints

Subject to:
2x + 3.2y <= 160
4x + 2.0y <= 240
y <= 40

Solving above contraint equation by Graphical method

2x +3.2 y =160

Taking X =0 , hence y =160/3.2=50

Putting y =0 , x = 160/2=80

(x,y ) = (0,50)(80,0)...(.1 )

Now 4x +2y =240

Putting x =0 , y =240/2 =120

Putting y = 0 , X =240/4=60

(X,y ) =(0,120) , (60,0)....(2)

And Y = 40........(3)

Plotting above equation

See image attached

y 1 120 DO 1980 60 A ex+2y7240 40 24=160 2 > 20 40 60 80 loo 120 X Value 40 2x + 3. 2 Finding eef intersecting intersecting p

a) Maximum quantity of X =16 and y =40

b)Maximum value of Z =$40.8